OpenStudy (mony01):

Can someone check if I did this right? I am trying to differentiate (sqrtx)cos(sqrtx)...

OpenStudy (mony01):

\[f'(x)=\sqrt{x}-\sin \sqrt{x}(\frac{ 1 }{ 2\sqrt{x} })+\frac{ 1 }{ 2\sqrt{x} }\cos \sqrt{x}\]

OpenStudy (primeralph):